# Gas price shocks, Uncertainty and Price setting: Evidences from Italian Firms

### v1

@@ 3 Empirical Analysis
3.1 Shock identification and impulse response functions
In order to identify a shock of nominal natural gas price, we propose the same framework as Kilian and Zhou (2022), but we apply it to the case of natural gas. We estimate a VAR with three variables: the real natural gas price, the log change in HICP, computed by excluding the price of natural gas, and firms year on year inflation expectations. The model is estimated at quarterly frequency from 1999Q4 to 2025Q2. Since the SIGE survey is conducted at the end of each quarter, all variables are measured at the same point in time—specifically, the final month of each quarter (March, June, September, December). This alignment preserves the within-month timing structure necessary for identification and ensures the proper alignment of the data.
Let $y_{t,m=3}=[\pi_{t,m=3}^{gas},\pi_{t,m=3}^{core},E_{t,m=3}\pi_{t+1}^{firms}]$ be a vector containing natural gas price $\pi_{t,m=3}^{gas}$, the log difference in HICP inflation excluding gas $\pi_{t,m=3}^{core}$ and year-on-year inflation expectations of firms $E_{t,m=3}\pi_{t+1}^{firms}$. The structural VAR model is given by:
$B_{0}y_{t,m=3}=\sum_{j=1}^{P}B_{j}y_{t-j,m=3}+w_{t,m=3}$ (1)
where $w_{t,m=3}$ is a vector of structural innovations. We can write the reduced form model as:
$y_{t,m=3}=\sum_{j=1}^{P}A_{j}y_{t-j,m=3}+u_{t,m=3}$ (2)
where $A_{j}=B_{0}^{-1}Bj$, j=1,2,,Pj=1,2,...,P, and the lag order $P$ is set to 4. We identify the nominal natural gas price shock by placing a combination of sign and zero restrictions on $B_{0}^{-1}$ by using the Bayesian approach of Arias, Rubio-Ramírez and Waggoner (2018) and adopting a uniform-Normal-inverse Wishart Prior. The nominal natural gas price shock is assumed to increase the real price of natural gas, as core inflation does to not respond in the same month as the price of natural gas. It increase headline inflation and inflation expectations of firms. The shock to core inflation reduces the real price of natural gas, since it does not respond immediately to an inflation shock, and increases core inflation and inflation expectations. Finally, a shock to firms’ inflation expectations has 0 effect on impact on real natural gas price and core inflation. The restrictions are reported in Equation 3
[…]
Figure 5 reports the impulse response functions to a one-standard-deviation natural gas price shock. The shock generates a sharp and persistent increase in the real natural gas price, with the effect gradually dissipating over approximately 10 quarters. The log change in headline inflation responds immediately, peaking on impact at around 15 basis points before returning to baseline within 5 quarters, consistent with the transitory nature of energy price shocks. In contrast, firms’ inflation expectations exhibit a more persistent response, with a positive response dying out in about 12 quarters.
Figure 5: Impulse response funcions of real natural gas price, headline inflation and firms’ inflation expectations to a natural gas price shock. The figure reports the impulse response functions of real natural gas price (left), headline inflation (center) and firms’ inflation expectations (right) to a natural gas price shock. Shaded areas represent 68% confidence bands. Sample: 1999:Q4–2025:Q2.
The impulse response functions show that natural gas price shocks are an important driver of firms’ inflation expectations. Figure 6 presents the historical decomposition of these expectations and of the log change in HICP excluding gas, isolating the contribution of each identified shock. Natural gas price shocks emerge as a primary driver of variation in firms’ annual inflation expectations throughout the sample period. These shocks predominantly exerted downward pressure on expectations, with a notable exception during the post-COVID-19 period (2021-2023), when they contributed substantially to the upward revision in firms’ inflation forecasts. This is unsurprising, given the high exposure of the Italian economy to natural gas supplies as its main energy source. Additionally, the appear to drive upward the log change in HICP during the same period.
Figure 6: Historical decomposition of firms’ year-on-year inflation expectations and log change in HICP excluding natural gas. The figure reports the historical decomposition of firms’ annual inflation expectations (top) and log change in HICP excluding natural gas price (bottom). Sample: 1999:Q4–2025:Q2.
3.2 Natural gas price shocks and firms’ pricing decisions
In order to evaluate how natural gas price innovations affect the pricing decisions of firms, we extract the identified structural shock from the first VAR. Then, we estimate a larger model, in which the shock is ordered as the first variable. Apart from natural gas price innovations, the model contains our estimated index of inflation uncertainty, the first moment of firms’ annual rate of change in their own prices and in their year-on-year expectations for their own prices, the annual HICP, unemployment rate, the short term interest rate for the Euro area and an industrial confidence index. The model is estimated by using the most advanced Bayesian techniques, adopting the hierarchical approach of Giannone, Lenza and Primiceri (2015) for the estimation of the hyperparameters and Lenza and Primiceri (2022) correction in order to account for the huge volatility of shocks during Covid. As all our variables are stationary, we estimate the model by shrinking the autoregressive term to 0. Our sample spans from the last quarter of 2002 until the second quarter of 2025. Due to the short sample and the large number of variables in the model, we estimate it with 2 lags. However, increasing the number of lags to 4 delivers nearly identical results.
@@ 3.2 Natural gas price shocks and firms’ pricing decisions
Figure 7: Impulse response functions of aggregate macroeconomic variables to a natural gas price shock. The figure reports the impulse response functions of inflation uncertainty undex, firms’ expected and realized annual price change, annual HICP, unemployment rate, short term interest rate and industrial confindence index to a natural gas price shock. Shaded areas represent 68% confidence intervals. Sample: 2002:Q4–2025:Q2.
Figure 7 reports the generalized impulse response functions to a one standard deviation natural gas price innovation, obtained by imposing recursivity in the model. As a consequence of the shock, annual inflation increases by about 10%. The effect is persistent and vanishes after roughly six quarters. Firms react by raising both their own prices and their expected prices in the short run, pointing to a rapid pass-through of higher energy costs. These responses peak within the first two quarters and then gradually fade away as weaker demand conditions and tighter monetary policy curb further price increases. At the same time, inflation uncertainty rises, suggesting that the shock not only pushes up the level of prices but also complicates firms’ forecasting. The industrial confidence index shows a short-lived improvement in the immediate aftermath of the shock, possibly reflecting temporary gains for some firms able to pass on costs, but it turns negative after the fifth quarter as higher energy prices depress demand and profitability. Overall, the adjustment remains temporary: the cost-push nature of the shock dominates in the short run, but the inflationary pressures dissipate once the economy absorbs the shock.
3.3 State dependent local projections
In order to evaluate the presence of non-linearities in firms’ price setting behavior in response to a shock to the price of natural gas, we employ state dependent local projections (Ramey and Zubairy, 2018). We determine states of high and low firms’ inflation uncertainty using the index constructed using the firm level microdata from the SIGE survey by adopting the approach of Binder (2017). Following Falck, Hoffmann and Hürtgen (2021) we scale the uncertainy index for the average inflation expectations of firms, in order to account for high inflationary periods, and we smooth it by applying a 4 period backward looking weighted moving average filter. We adopt two different rules for determining states of high inflation uncertainty. According to the first rule, the transition between states of high and low probabilities of inflation uncertainty is governed a logistic function $Z(\hat{\Delta}_{t-1})\in[0,1]$ of the lagged index, defined as:
@@ 3.3 State dependent local projections
where $\hat{\Delta}_{t}$ is the state variable, $\eta$ is a parameter that determines the steepness of the transition, and $\mu$ and $\sigma_{\hat{\Delta}}$ represent the median and standard deviation of the state variable, respectively. Following Falck, Hoffmann and Hürtgen (2021), we set $\eta=5$ and we compare probability regimes with the one obtained by adopting a simple threshold rule: the economy is in the high-uncertainty state $Z(\cdot)$ with probability 1 whenever $\hat{\Delta}_{t-1}>0$, and 0 otherwise.
Figure 8: Inflation uncertainty states. The figure displays the estimated probability of the high inflation uncertainty state, based on the measure by Binder (2017), and as defined by the smooth transition logistic function (light gray areas) and the threshold rule (dark gray areas). Sample: 2004:Q3–2025:Q2.
@@
The transition probabilities are reported in Figure 8, together with the smoothed and scaled index of inflation uncertainty. The probability of being in the high uncertainty state increases in the aftermarth of the GFC and remains high during the sovereign debt crisis. It falls to 0 after 2019, and then it surges again in the aftermarth of the Covid-19 episode and during the European energy crisis, remaining high until the end of our sample. We then estimate the following set of state dependent local projections:
@@
$x_{t+h}=$ $\bigg[\tilde{\beta}^{Z}_{x}(h)w^{gas}_{t}+\textit{controls}(x,Z)\bigg]Z(\hat{\Delta}_{t-1})$ $+\bigg[\tilde{\beta}^{\bar{Z}}_{x}(h)w^{gas}_{t}+\textit{controls}(x,\bar{Z})\bigg]\bar{Z}(\hat{\Delta}_{t-1})+\epsilon_{t+h}$ (5)
Where $\bar{Z}\equiv 1-Z(\hat{\Delta}_{t-1})$. The estimated impulse response function are reported in Figure 9. The top row shows the impulse responses obtained by adopting the simple threshold state transition rule, while the responses at the bottom are obtained by adopting the smooth transition rule based on the logistic function of Equation 4. Strong nonlinear effects emerge in the transmission of natural gas price shocks to firms’ price-setting decisions. Under the high uncertainty scenario, firms are confident that they will be able to pass the increased costs to their buyers. They increase their own prices persistently and raise their expectations for next year’s prices. However, this confidence appears to be short-lived. As the recession hits, price increases stop. Under the low uncertainty scenario, by contrast, firms anticipate the recessionary impact of the shock. They revise their prices downward and adjust their expectations accordingly.
Threshold rule
Smooth transition rule
@@
Figure 9: Firms expected and realized annual price change state-dependent response to a natural gas price shock. Estimation follows the lag-augmentation method of Montiel Olea and Plagborg‐Møller (2021), with 4 lags. The responses in the top row are estimated using a clear threshold rule. The responses at the bottom row are estimated using the smooth transition probability rule. Shaded bands represent 68% confidence intervals based on Newey-West standard errors. Sample: 2004Q3–2025Q2.

### v3

@@ 3 Empirical Analysis
3.1 Identification through external instrument
Instrument construction
In order to identify a supply shock to natural gas, we follow the approach of kanzig2023unequal (kanzig2021macroeconomic, kanzig2023unequal). As alessandri2025natural, we exploit news extracted from Refinitiv. Differently from them, we focus on a smaller number of news from Reuters about interruptions of gas supplies to Europe which are highly likely to be exogenous to the economy and unexpected by the markets. We include episodes of unexpected interruptions caused by extreme weather event, like the cold snap that hit Russia in 2012 forcing it to reduce the export of natural gas to satisfy its domestic demand, and interruptions in the supply caused by terrorist attacks, protests and unplanned maintenance in Libya and Algeria. Furthermore, we include incidents like the explosion that hit the Baumgarten natural gas facility in Austria in 2017 causing the interruption of the flow of gas to the Italian Hub of Tarvisio and the fire in the Freeport liquefied natural gas facility in US which reduced the supply of LNG to Europe. Additionally, for the 2022 period we include some announcements from Russia which have been perceived as a surprise from the market and which contain news about future reduction in the supply of natural gas. As an example, the announcement by Putin that Russia was going to stop the deliveries of natural gas, unless the payments were made in robles in March 2022 concretely reduced the amount of purchasable commodity for countries which could not or did not want to comply with the request, and has been followed by a positive movement of the Dutch TTF (alessandri2025natural). We avoid to include the sabotage of the Nord Stream 2 pipeline since the latter never actually delivered natural gas to Europe, thus the event did not concretely change the amount of commodity supplied to European countries. We also exclude the strikes in gas production facilities in Norway during 2022, since these episodes were motivated by the request of Norvegian workers to increase wages in response to inflationary pressures. Furthermore, to reduce the contamination of the instrument with geopolitical shocks we do not include the Russian invasion of Ukraine in February 2024 even though it has been followed by a massive surge in the Dutch TTF.
Table 1: Natural gas surprises
Event date Event description
1 January 2006 Russia stops flows to Ukraine due a payment dispute, blocking gas flows to Italy and Austria
22 February 2011 Due to protests and civil war, the Greenstream pipeline stopped delivering gas from Libya to Italy
31 January 2012 An unexpected cold snap forces Russia to reduce the supply of gas to Europe to satisfy its domestic demand
17 January 2013 Supplies from Algeria to Italy fall because of a terrorist attack in Almenas Gas Field
4 March 2013 Clashes with militias in Mellitha complex in Libya completely stops the gas flows to Italy.
30 September 2013 Protests in Libya cause the shut down of a gas pumping station, halting the gas flows to Italy
23 January 2014 Gas flows from Algeria to Italy stopped due to quality issues. Received gas had to be sent back
29 April 2015 Libya shuts down a gas field in the east due to protests
13 December 2017 A blast in Austrian hub of Baumgarten interrupts gas transit to Tarvisio hub in Italy
05 February 2022 An unplanned maintenance at Mellitah complex considerably reduces gas flows from Libya to Italy
23 March 2022 Putin asks European countries to pay gas deliveries in Rubles otherwise gas flows will stop
8 June 2022 Fire in the Freeport LNG facility in the US
14 June 2022 NS1 operate at limited capacity due to turbines stuck in Canada
31 August 2022 Flows of gas from NS1 shut down permanently
2 October 2022 Russia halts gas deliveries to Italy trough the TAG pipeline
  • Notes: The table reports the natural gas surprises we use to construct the instrument.
We end up isolating 15 surprises which are reported in Table 1. We construct a daily instrument which takes value of 1 in correspondence of any of the surprises. Then, we aggregate the instrument at the monthly level.
IV-VAR specification
We specify a monthly VAR including six variables. We take the day ahead Dutch TTF as a measure of the spot price of natural gas. Then, in the spirit of adolfsen2024gas, we include the Italian gas quantities, gas inventories and industrial production. We also control for the world industrial production and for the Prezzo Unico Nazionale (PUN) as a measure of the Italian spot electricity price. [^6] Data about gas quantities and gas inventories are provided by the Italian Ministero dell’ambiente e della sicurezza energetica. Gas quantities are defined as national production, plus imports minus exports. Gas inventories are given by the cumulative sum of the variation in inventories, given a base value of 7000 millions of standard cubic meters. Both gas quantities and gas inventories are seasonally adjusted by using the X-13 ARIMA seasonal adjustment method. Data about the PUN are provided by Gestore Mercati Energetici. World industrial production is provided by BH2019. Monitoring the response of the PUN to a gas supply shock is informative because, in a uniform-price electricity market with gas-fired units often setting the marginal cost, like the Italian one, changes in the TTF day-ahead price feed into generators’ bids and hence the Italian spot price with short lags; the PUN therefore reflects downstream pass-through from gas to electricity price. Controlling for world industrial production purges the gas-price equation of global demand pressure, sharpening identification of a gas supply shock. The VAR is estimated in log levels with a constant from 2004M1 until 2025M6. The lag order is set at 12. We assess instrument relevance by performing an F-test in the regression of the TTF residuals from the VAR on the instrument, following olea2021inference. An F-statistic above 10 indicates a strong instrument. In our case, the F-statistic equals 13.6, confirming high instrument strength.
Figure 5: Gas supply shock. The figure shows the natural gas supply shock obtained from the IV-VAR (left) and the corresponding correlogram (right). Sample: 2005:M1–2025:m6.
We also conduct the F-test allowing for heteroskedasticity, which delivers an F-statistic of 15.18.777The instrument explains around 4.9% of the TTF residual. Figure 5 reports the structural shock extracted from the IV-VAR and the corresponding correlogram.
Shock exogeneity
We test the exogeneity of our shock with respect to oil and energy-related shocks identified in the literature using a Wald test of joint significance. Specifically, we regress our shock measure on the contemporaneous and 12 lagged values of each structural innovation considered, as well as on 12 of its own lags. We consider the oil price news shock of Kanzig2020, the carbon price shock of kanzig2023unequal, the oil demand shocks of CALDARA20191 and BH2019, the oil supply,inventories and demand shock of BH2019, the oil price expectation shock of Baumeister2023, and the oil supply shocks of CALDARA20191, kilian2009not and kilian2008exogenous. Results are reported in Table 2. The Wald tests do not reject the null of no contamination for any proxy, supporting their exogeneity. Only the BH2019 oil inventories shock approaches significance at the 10% level.
Table 2: Wald test results.
Innovation Sample Wald statistic P-value
Kanzig2020 Oil price news 2000Q4 - 2024Q2 7.8868 0.7939
kanzig2023unequal Carbon price 2000Q4 - 2019Q4 16.0719 0.188
CALDARA20191 Oil demand 2000Q4 - 2015Q4 12.0753 0.4396
CALDARA20191 Oil supply 2000Q4 - 2015Q4 10.8998 0.5375
BH2019 Oil supply 2000Q4 - 2025Q2 9.2402 0.6823
BH2019 Oil consumption demand 2000Q4 - 2025Q2 11.9281 0.4515
BH2019 Oil inventories 2000Q4 - 2025Q2 20.1281 0.0647
Baumeister2023 Oil price expectation 2000Q4 - 2023Q1 6.1135 0.9102
kilian2009not Oil supply 2000Q4 - 2009Q4 5.7186 0.9296
kilian2008exogenous Oil Supply 2000Q4 - 2016Q4 8.2033 0.7690
  • Notes: The table reports Wald test statistics for the joint significance of coefficients. The null hypothesis is that all coefficients are zero, indicating no contamination.
Energy Market Responses to a Gas Supply Shock
Figure 6 shows the impulse response functions from the VAR estimated by using the external instrument approach. As in kanzig2021macroeconomic, we can interpret the impulse response functions as elasticities, as all the variables in the VAR are in logs. We normalize the shock to increase the gas price, measured by the TTF, by 1%. We can see that price variables immediately react to the shock. The TTF responds positively and the response reverts to 0 in a bit more than an year. The increase in the price of natural gas is then transmitted to the national electricity price, measured by the PUN, which increases strongly, peaking within a few months and then slowly reverting. This confirms pass-through from gas to electricity: gas-fired plants set the marginal price, so the electricity market mirrors gas dynamics with a short lag. About half of the increase in the gas price is transmitted to the electricity price on impact. After a few months, gas inventories decrease of about 0.1%. Higher gas prices coincide with withdrawals from storage before restocking once prices ease. Gas quantity initially do not respond to the shock. After 10 months, we observe a small drop in gas quantity followed by a gradual recovery. Industrial production is not affected by the shock, consistently with its transitory nature.
Gas supply shock
[Uncaptioned image] Figure 6: Impulse response functions of energy market variables to a natural gas supply shock. The figure reports the impulse response functions of natural gas price, gas inventories, gas quantity, industrial production, PUN and world industrial production to a natural gas supply shock. Shaded areas represent 68% and 90% confidence intervals, computed using a moving block bootstrap procedure. Source: Ministero dell’ambiente e della sicurezza energetica, Gestore Mercati Energetici, ISTAT, Refinitiv Workspace, BH2019. Sample: 2002:Q4–2025:Q2.
3.2 Natural gas supply shocks and firms’ behavior
Bayesian VAR
In order to evaluate how natural gas supply innovations affect the pricing decisions of firms, we estimate model shown in Section 3.1 and we extract the identified structural shock. Then, we estimate a proxy-internalizing VAR, placing the shock of interest as the first variable in the system and imposing recursivity (BRULU2022). [^8] According to PMW2021, the advantage of this approach over the equivalent SVAR-IV method of SW2018, besides its simplicity, is that the estimated impulse responses are valid even if the shock of interest is non-invertible. Apart from our estimated innovation, the model contains our index of inflation uncertainty, the average firms’ annual rate of change in their own prices and the average expected change in their prices over the next year, the annual rate of change in their assessment about order-book levels and about the stock of finished products and their inflation expectations. Firms expectations from SIGE are measured at quarterly frequency, at the final month of each quarter. In order to preserve the proper time alignment, we follow AB2023 and we take the end of the quarter values of the gas supply shock. Our sample spans from 2005Q4 until 2025Q2.
Gas supply shock
@@ 3.2 Natural gas supply shocks and firms’ behavior
[Uncaptioned image] Figure 7: Impulse response functions of firms’ related variables to a natural gas supply shock. The figure reports the impulse response functions of inflation uncertainty index, firms’ expected and realized annual price change, firms’ assessed order-book levels and stock of finished products and annual inflation expectations to a natural gas supply shock. Shaded areas represent 68% confidence intervals. Source: Bank of Italy, Survey on inflation and growth expectations, EUROSTAT, European Commission - ECFIN. Sample: 2005:Q4–2025:Q2.
Due to the limitations of our sample size, we resort to Bayesian methods for recovering the impulse response functions. We adopt the hierarchical approach of giannone2015prior for the optimization of the hyperparameters and the lenza2022estimate correction, in order to account for the huge volatility of shocks during Covid. As all our variables are stationary, we estimate the model by shrinking the autoregressive term to 0. We set the number of lags to 4. Figure 7 reports the generalized impulse response functions to a shock that increases the TTF by 1%. All dependent variables in this specification are expressed in annual percentage changes. The results are consistent with a cost-push disturbance. On impact, the annual rate of change in firms’ prices increases by 5% and the effect fades in a bit more than one year. Expectations about future prices display a similar but shorter-lived pattern. The year on year change jumps higher with respect to price realizations on impact, and reverts to 0 in half an year. Inflation uncertainty increases sharply for about one year, suggesting that the shock not only pushes up the level of prices but also complicates firms’ forecasting. Inflation expectations move in the same direction and with a similar scale as price expectations, but the increase in their annual growth rate is as persistent as the one for firms’ realized prices.
State dependent local projections
adolfsen2024gas argue that the pass-trough of a natural gas supply shock in the Euro Area is characterized by a non linearity in the level of tightness in the labor market, with effects that are considerably magnified when the economy operates near full capacity. We show that the pass-through of natural gas supply shocks is also characterized by substantial non-linearities related firms’ ability to accurately forecast inflation. When inflation uncertainty increases, it becomes more difficult for firms to form an expectation about the future level of prices. This affects the way in which they respond to the shock of the price of natural gas. Under high uncertainty scenarios, the pass through is larger and firms persistently increase their prices. When instead uncertainty about future inflation is low they reduce their prices in order to preserve their market shares, anticipating the recessionary effects of the shock. We illustrate this by adopting state dependent local projections (ramey2018government). We determine states of high and low firms’ inflation uncertainty using the index constructed using the firm level micro data from the SIGE survey by adopting the approach of binder2017measuring. Following F2021 we scale the uncertainty index for the average inflation expectations of firms, in order to account for high inflationary periods, and we smooth it by applying a 4 period backward looking weighted moving average filter. We assume that the transition between states of high and low probabilities of inflation uncertainty is governed a logistic function $Z(\hat{\Delta}_{t-1})\in[0,1]$ of the lagged index, defined as:
@@
where $\hat{\Delta}_{t}$ is the state variable, $\eta$ is a parameter that determines the steepness of the transition, and $\mu$ and $\sigma_{\hat{\Delta}}$ represent the median and standard deviation of the state variable, respectively. Following F2021, we set $\eta=5$. [^9] Results are consistent across different values for parameter $\eta$. In order to account for the non linearities in unemployment shown by adolfsen2024gas, we do the same for the unemployment rate, determining the probabilities of having a tight labor market.
Figure 8: Inflation uncertainty and labor market states. Top panel: estimated probability of high inflation uncertainty following binder2017measuring, computed via smooth transition logistic function (shaded areas). Bottom panel: estimated probability of tight labor market based on unemployment, computed via smooth transition logistic function (shaded areas). Blue lines represent the underlying state variables. Source: Bank of Italy, Survey on inflation and growth expectations, EUROSTAT. Sample: 2007:Q1–2025:Q2.
@@
The transition probabilities are reported in Figure 8, together with the smoothed and scaled index of inflation uncertainty and the unemployment rate. The probability of being in the high uncertainty state increases in the aftermath of the GFC and remains high during the sovereign debt crisis. It falls to 0 after 2019, and then it surges again in the aftermarth of the Covid-19 episode and during the European energy crisis, remaining high until the end of our sample. Instead, the probabilities of having a tight labor market are high until 2011. Then they remain at 0 until 2021, with the exception of a single episode, coinciding with the onset of the COVID pandemic in 2020. We then estimate the following set of state dependent local projections:
@@
$x_{t+h}=$ $\bigg[\tilde{\beta}^{Z}_{x}(h)w^{gas}_{t}+\textit{controls}(x,Z)\bigg]Z(\hat{\Delta}_{t-1})$ $+\bigg[\tilde{\beta}^{\bar{Z}}_{x}(h)w^{gas}_{t}+\textit{controls}(x,\bar{Z})\bigg]\bar{Z}(\hat{\Delta}_{t-1})+\gamma_{D}D_{t}+\epsilon_{t+h}$ (3)
Where $Z(\hat{\Delta}_{t-1})$ indicate the probability of being in the state of high inflation uncertainty, $\bar{Z}\equiv 1-Z(\hat{\Delta}_{t-1})$, $\textit{controls}(x,\bar{S})$ is a set of control variables including the interaction between $w^{gas}_{t}$ and the probability of being in the tight labor market state $S(\hat{\Delta}_{t-1})$ and lags of the shock and of the dependent variable $x_{t}$. [^10] In our baseline specification, we include two lags of the dependent variable and two lags of the shock to preserve degrees of freedom. Results are robust to specifications with four lags of both variables. $D_{t}$ is a dummy variable that assumes value of 1 in correspondence of 2020Q2 and then decays exponentially over the subsequent quarters. The estimated impulse response function are reported in Figure 9.
Gas supply shock
@@
Figure 9: Firms annual price change state-dependent response to a gas supply shock. The figure shows the state dependent response of firms’ prices to a gas supply shock. Estimation follows the lag-augmentation method of MOPM2021. The number of lags of the dependent variable, the control variables and the shock in each state are determined by minimizing the AIC. Shaded bands represent 68% confidence intervals based on Newey–West standard errors. Source: Bank of Italy, Survey on inflation and growth expectations, EUROSTAT. Sample: 2007Q1–2025Q2.
When inflation uncertainty is high, firms respond to a natural gas supply shock by persistently raising their prices. This behavior reflects precautionary pricing motives: firms face higher perceived risks about future input costs and the duration of the energy shock, and therefore widen their mark-ups to protect margins against potential future cost increases. The widespread and salient nature of energy costs also facilitates price coordination across firms, reducing competitive pressure and allowing a stronger pass-through of cost shocks. Under low uncertainty instead, firms’ prices fall in response to the shock to the gas supply. The negative effect of the shock on aggregate demand is clear. Firms anticipate the recession and do not increase their prices in order to preserve their market share.